solve for x. sqrt 3x=9.
Answer by Maths68(1474)

sqrt 3x=9
Taking square on both sides
(sqrt 3x)^2=(9)^2
3x=81
divide by 3 both sides of the above equation
3x/3=81/3
x=27
Check:
sqrt 3x=9
put the value of x in above equation
sqrt 3(27)=9
sqrt 81=9
sqrt (9*9)=9
9=9